OpenStudy (anonymous):

Solve: 2(x + 2) − (x + 1) = 8

OpenStudy (aravindg):

use ditributive property

OpenStudy (anonymous):

First, distribute the 2 to the first parentheses, and then the (-) to the second set of parentheses. What do you get?

OpenStudy (aravindg):

*distributive

OpenStudy (aravindg):

\[\large \boxed{a \times (b +c)=ab+ac }\]